Add 27 to the measure of an angle and the result is half of the angles supplement what is the measure of the angle
Answer by Boreal(15235)

angle is x
supplement is 180-x
x+27=(1/2)(180-x)
x+27=90-(x/2)
(3x/2)=63
x=42 deg ANSWER
supplement is 138 deg
Half the supplement is 69 deg, which is 27 more than 42.